A cleaning solution was made by mixing 10 qt of a 4% baking soda solution and 8 qt of a 45% soap solution. Find the concentration of the cleaning solution.
Answer by alicealc(293)

baking soda = 4/100 * 10 qt = 40/100 qt = 0.4 qt
soap = 45/100 * 8 qt = 360/100 qt = 3.6 qt

concentration
= (0.4+3.6)/(10+8) * 100%
= 4/18 * 100%
= 400/18 %
= 22.22 %